OpenStudy (anonymous):

find an equation of the line that passes through the given point and has the indicated slope m. question 1: (0,-2), m = 3 question 2: (-3, 6), m = -2 question 3: (4,0), m = -1/3

OpenStudy (nincompoop):

use the point-slope form \[y-y_1 = m(x-x_1) \] then solve for y

OpenStudy (anonymous):

so what you want to do is plug the slope value (m) into the basic y=mx+b equation. then plug the point (x,y) into the equation and solve for b

OpenStudy (nincompoop):

an example given a point (4, 1) and a slope of 5, find the equation of a line solution: first assign \[m = 5\] \[x_1 = 4\] \[y_1 = 1\] using the point-slope form \[y - 1 = 5 (x-4) \rightarrow y-1 = 5x - 20\] solve for y \[y - 1 + 1 = 5x - 20 + 1 \rightarrow y = 5x -19\]